### v2

# MobileRL: Online Agentic Reinforcement Learning for Mobile GUI Agents

Abstract
Building general-purpose graphical user interface (GUI) agents has become increasingly promising with the progress in vision language models. However, developing effective mobile GUI agents with reinforcement learning (RL) remains challenging due to the heavy-tailed distribution of task difficulty and the inefficiency of large-scale environment sampling. We present an online agentic reinforcement learning framework MobileRL to enhance GUI agents in mobile environments. Its core component is the Difficulty-ADAptive GRPO (AdaGRPO) algorithm. In AdaGRPO, we design difficulty-adaptive positive replay and failure curriculum filtering to adapt the model to different task difficulties. We introduce the shortest-path reward adjustment strategy to reshape rewards concerning the task length in multi-turn agentic tasks. Those strategies jointly stabilize RL training, improve sample efficiency, and generate strong performance across diverse mobile apps and tasks. We apply MobileRL to two open models (Qwen2.5-VL-7B-Instruct and GLM-4.1V-9B-Base). The resultant MobileRL-9B model achieves state-of-the-art results in terms of success rates on both AndroidWorld (80.2%) and AndroidLab (53.6%). The MobileRL framework is open-sourced at https://github.com/THUDM/MobileRL.
11footnotetext: YX and XL contributed equally.22footnotetext: Work partially done while these authors interned at Z.AI.
Figure 1: Left and Center: Task success rates on AndroidWorld (Rawles et al., 2024) and AndroidLab (Xu et al., 2024); hatched areas indicate gains from MobileRL on top of the SFT model. Right: Trajectory-level success rate curves on AndroidWorld train and test sets during RL training.
1 Introduction
GUI agents—powered by vision language models—have enabled zero-shot interaction with web pages and mobile interfaces (Hong et al., 2023; OpenAI, ; Bai et al., 2025; Liu et al., 2024). To improve them, significant efforts have focused on supervised fine-tuning or offline imitation learning over static expert demonstrations (Rawles et al., 2023; Xu et al., 2024; Bai et al., 2024; Lu et al., 2025). However, these methods suffer from limited behavior coverage and poor error recovery (Chang et al., 2022). Reinforcement learning (RL) with verifiable rewards presents a promising alternative (DeepSeek-AI et al., 2025; Hou et al., 2025). Yet, existing datasets with single-step expert labels  (Qin et al., 2025; Luo et al., 2025) are insufficient for training or evaluating policies on agentic tasks (i.e., planning and reasoning over multi-step action sequences). Although early progress has been made on online learning for GUI agents (Bai et al., 2024; Dong et al., 2025; Dai et al., 2025), efficiently scaling agentic RL in interactive mobile simulators remains largely unexplored.
Specifically, it faces the following technical challenges: (i) Complex instruction following under sparse positive signals: base models usually struggle to reliably produce correct action commands for complex, GUI-specific instructions. Due to the heavy cost and latency of mobile emulation, correctly-executed rollouts are rare, resulting in data-inefficient early exploration. (ii) Large and unstable task difficulty spectrum: some tasks can succeed with multiple rollouts, while others are persistently unsolvable for the model. Naive sampling wastes computational budget and under-utilizes scarce but informative trajectories (Xu et al., 2024). (iii) Sampling bottlenecks in large-scale mobile environments: deploying and managing hundreds of concurrent mobile instances is resource-intensive and hard to reproduce across setups. Low sampling throughput further limits both the scale and efficiency of online agentic RL.
To address these challenges, we present an adaptive online agentic RL framework MobileRL for advancing mobile GUI agents. MobileRL consists of three components: reasoning-free supervised fine-tuning (SFT), reasoning SFT, and agentic RL. The two SFT stages provide a warm-up for RL. Specifically, reasoning SFT enhances the handling of long and compositional instructions, reduces costly on-policy trials in mobile simulators, and enables the broad use of open or human-labeled datasets without relying on proprietary models.
To enable effective online agentic RL, we introduce Difficulty–ADAptive Group Relative Policy Optimization (AdaGRPO). Built upon group relative policy optimization (GRPO) (Shao et al., 2024), its core idea is to adapt optimization to instance difficulty and explicitly reward solution efficiency. AdaGRPO designs three key strategies: (i) Difficulty-Adaptive Positive Replay (AdaPR) maintains a curated buffer of challenging, high-quality trajectories and balances them with on-policy samples (Mnih et al., 2015; Zha et al., 2019). In sparse-reward mobile environments, difficult successes are rare yet highly informative; replaying them amplifies their learning signal and stabilizes policy updates. (ii) Failure Curriculum Filtering (FCF), as a simplified version of curriculum learning (Matiisen et al., 2019; Narvekar et al., 2020), down-weights persistently unsolvable tasks using online difficulty statistics, reallocating computational budget toward feasible instances. Given the heavy-tailed difficulty distribution observed in mobile agent benchmarks (Xu et al., 2024; Rawles et al., 2024), pruning hard dead-ends improves sample efficiency while retaining signal from recoverable failures. (iii) Shortest‐Path Reward Adjustment (SPA) reshapes the reward function based on completion length, assigning higher returns to shorter solutions. Length-sensitive rewards counteract bias toward verbose and better align with user preferences in mobile interaction contexts.
We implement MobileRL in a Verl-based framework (Sheng et al., 2024), which supports multi-task, multi-turn agentic RL training. Unlike previous Android simulator implementations (Toyama et al., 2021; Rawles et al., 2024)—which generally do not support true concurrent execution, our framework sustains high throughput that orchestrates hundreds of Dockerized Android virtual devices (AVDs) across multiple machines. This setup enables concurrent interaction with over 1,000 environments while preserving reproducibility. Since most open-source benchmarks and simulators are built upon the Android operating system (Toyama et al., 2021; Rawles et al., 2024), this design ensures seamless compatibility and faithful reproduction of environment behaviors.
We train MobileRL on Qwen2.5-VL-7B-Instruct (Bai et al., 2025) and GLM-4.1V-9B-Base (Team et al., 2025), producing MobileRL-7B and MobileRL-9B, respectively. MobileRL-9B lifts the success rates to 80.2%  on AndroidWorld and 53.6%  on AndroidLab, significantly outperforming previous state-of-the-art results (64.2% and 41.2%, respectively). MobileRL-7B outperforms the much larger 72B-parameter models UI-Tars-1.5 (Qin et al., 2025) and UI-Genie-Agent (Xiao et al., 2025) (e.g., +16% on AndroidWorld), despite being substantially smaller in scale. Also, extensive ablation studies demonstrate the effectiveness in the design of AdaGRPO.
In summary, our contributions are as follows:
  • MobileRL Framework & scalable sampling: We develop MobileRL with a two-stage warm-up followed by online agentic RL for mobile GUI agents. We further establish a distributed sampling implementation that coordinates hundreds of Dockerized Android virtual devices (AVDs), enabling reproducible large-scale training on Android benchmarks.
  • AdaGRPO Algorithm: We introduce Difficulty–Adaptive Group Relative Policy Optimization (AdaGRPO), which extends GRPO with (i) AdaPR for replaying challenging successful trajectories, (ii) FCF for down-weighting persistently unsolved tasks, and (iii) SPA for length-sensitive reward shaping, thereby accounting for instance difficulty and solution efficiency.
  • Empirical Results: Training on Qwen2.5-VL-7B-Instruct and GLM-4.1V-9B-Base yields MobileRL-7B and MobileRL-9B. MobileRL-9B reaches 80.2% on AndroidWorld and 53.6% on AndroidLab, surpassing previous bests (64.2% / 41.2%).
2 MobileRL
Figure 2: Overview of MobileRL. It consists of 1) reasoning warm-up with both reasoning-free SFT and reasoning SFT and 2) online agentic RL with AdaGRPO. In AdaGRPO, the warmed-up policy interacts with mobile environments to generate rollouts, which are scored by shortest-path reward adjustment (SPA). High-quality positive trajectories are stored in the AdaPR buffer, while low-performing rollouts are pruned via failure curriculum filtering.
We study mobile GUI agents and introduce the MobileRL framework, as shown in Figure 2, which aims to address three key challenges in interactive mobile environments: (i) following complex instructions under sparse and delayed rewards; (ii) handling a heavy–tailed and unstable task difficulty distribution; and (iii) overcoming large–scale sampling bottlenecks in mobile emulators.
Given a natural–language instruction (e.g., “open the calendar and add an event for tomorrow at 3 pm”), the agent autonomously performs closed–loop interactions with the mobile device. First, it perceives the current screen, grounds UI elements, and executes a sequence of actions without human intervention. The feedback is sparse and it can be observed only upon successful task completion, at which point the interaction terminates or a predefined horizon is reached. The goal is to learn a policy that generates strong performance across applications and tasks, minimizes unnecessary interactions, and maximizes task success.
Problem Formulation.
We model the mobile GUI agent as a finite-horizon Markov Decision Process (MDP) (Littman, 2009) $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,H,\mu_{0})$. The state space $\mathcal{S}$ contains all possible GUI states; a concrete state at time $t$ is $s_{t}\in\mathcal{S}$, which comprises a screenshot of the device screen together with the structured UI hierarchy parsed from XML metadata (Cf. Appendix C). The action space $\mathcal{A}$ is a finite set of atomic GUI operations; an action at time $t$ is $a_{t}\in\mathcal{A}$, including primitives such as Tap, Swipe, Type, Long Press, Launch, Home, Back, Wait, and a terminating action Finish. $P(s_{t+1}\mid s_{t},a_{t})$ represents the stochastic transition mechanism of the Android OS and installed applications, capturing the uncertainty over possible next states. We consider a finite horizon $H\in\mathbb{N}$. The initial condition is drawn from a joint distribution $\mu_{0}$ over instruction–state pairs, i.e., $(s_{0},c)\sim\mu_{0}$. Here, $c$ is the natural-language task instruction provided once at the beginning of the episode (t=0t{=}0). At each timestep $t=0,1,\dots$, the agent observes only the current state $s_{t}$ and samples an action according to a policy $\pi_{\theta}$, i.e., $a_{t}\sim\pi_{\theta}(\cdot\mid s_{t},c)$. The environment then transitions to $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$. An episode yields a trajectory $\tau=\big((s_{0},a_{0}),(s_{1},a_{1}),\ldots,(s_{T},a_{T})\big)$ and terminates either when the agent intentionally selects Finish in a success state or when the horizon is reached, with $T\leq H$. The reward is assigned only after task completion, such that $R(\tau)=r(s_{T},a_{T})$ with $r(s_{T},a_{T})\in\{0,1\}$ indicating success (1) or failure (0). Consequently, learning maximizes the success probability: $\theta^{\star}=\arg\max_{\theta}\;\mathbb{E}_{(s_{0},c)\sim\mu_{0},\;\tau\sim\pi_{\theta}}\!\bigl[R(\tau)\bigr]$.
2.1 The MobileRL Framework
To build a powerful mobile use agent, we present the MobileRL framework. It comprises three components: reasoning-free supervised fine-tuning (SFT) on expert demonstration data, an iterative warm-up stage by reasoning SFT, and agentic RL with a difficulty–adaptive policy optimization strategy we developed in this work.
Reasoning-Free SFT.
In agentic RL training, sampling in virtual-device environments is usually inefficient; thus, starting online RL directly from a base model was found to be excessively time-consuming. Therefore, we perform SFT with expert demonstration data obtained by following the data collection protocol of (Xu et al., 2024), supplemented with the training split of the publicly available AndroidControl dataset (Li et al., 2024). Note that this data is reasoning-free.
Reasoning SFT.
To further construct a stronger reasoning policy initializer, we perform reasoning SFT via an iterative reasoning refinement strategy over the expert dataset. Manually collected expert demonstrations dataset for mobile use often contains only the final action sequence, omitting intermediate reasoning. Training solely on such “black-box” trajectories yields opaque policies. We leverage an off-the-shelf instruction model to bootstrap a reasoning-augmented training set from raw demonstrations, yielding a structured and transparent policy initialization. Concretely, we iteratively build the reasoning instruction–tuning pairs in three stages:
  • Bootstrap sampling: For each task $x$ with expert answer $a^{*}$, the Instruct model $M$ generates diverse candidate reasoning–action pairs $(c_{k},a_{k})$. Whenever $a_{k}=a^{*}$, we retain $(x,c_{k},a^{*})$ in $\mathcal{D}_{R}$.
  • Supervised fine-tuning: Train an initial reasoning policy $\pi^{R}_{0}$ on $\mathcal{D}_{R}$.
  • Iterative refinement: At iteration $t$, $\pi^{R}_{t}$ proposes candidates; those matching $a^{*}$ are scored by correctness. The best explanation $c^{*}$ is added to $\mathcal{D}_{\mathrm{new}}$, and $\pi^{R}_{t+1}$ is obtained by fine-tuning.
The resulting reasoning-oriented fine-tuning corpus is trained for two epochs to produce the reasoning warm-up model used for agentic RL training.
Agentic RL.
During agentic RL (multi-turn) training, we face the challenges of immediate reward assignment and sampling efficiency, which are discussed and addressed by building upon the group relative policy optimization (GRPO) (Shao et al., 2024).
Briefly, GRPO advances proximal policy optimization (Schulman et al., 2017) by replacing the learned value baseline with an on-the-fly, group-relative baseline computed from a set of trajectories for the same task. Given a shared initial condition $(s_{0},c)\sim\mu_{0}$, we sample a group of $G$ trajectories $\mathcal{G}=\{\tau_{1},\ldots,\tau_{G}\}$ by rolling out $\pi_{\theta_{\mathrm{sample}}}$. Let $T_{i}$ denote the number of steps for trajectory $\tau_{i}$, $s_{i,t},a_{i,t}$ be the state and action at step $t$, and the trajectory-level reward for $\tau_{i}$ be $R(\tau_{i})\in\{0,1\}$. We define the group-relative trajectory-level advantage for any step $(s_{i,t},a_{i,t})$ on trajectory $\tau_{i}$ as $\hat{A}_{i,t}\;\equiv\;\frac{R(\tau_{i})-\mu_{R}}{\sigma_{R}}\,$ , where $\mu_{R}$ and $\sigma_{R}$ are the mean and standard deviation of $\{R(\tau_{j})\}_{j=1}^{G}$. The GRPO loss can be written in empirical form as
$\mathcal{L}^{\mathrm{GRPO}}(\theta)=-\,\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\min\!\Bigl(\rho_{i,t}(\theta)\,\hat{A}_{i,t},\;\mathrm{clip}\!\bigl(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon\bigr)\,\hat{A}_{i,t}\Bigr)$ (1)
where $\rho_{i,t}(\theta)=\frac{\pi_{\theta}(a_{i,t}\mid s_{i,t})}{\pi_{\theta_{\mathrm{sample}}}(a_{i,t}\mid s_{i,t})}$ is the token-wise importance sampling (IS) ratio. We add the KL loss $\beta\,D_{\mathrm{KL}}\!\bigl(\pi_{\theta}\,\|\,\pi_{\theta_{\mathrm{ref}}}\bigr)$ to prevent the model from deviating too much from the prior distribution.
2.2 Difficulty–Adaptive GRPO
We develop Difficulty–Adaptive GRPO (AdaGRPO) with three strategies—shortest-path reward adjustment (SPA), difficulty–adaptive positive replay (AdaPR), and failure curriculum filtering (FCF)—to address the challenges faced in mobile agentic RL.
First, in multi‐turn mobile agentic tasks—where immediate rewards are absent within a single round, unlike single‐turn settings (Lu et al., 2025; Team et al., 2025)—the reward allocation strategy must be redesigned. Beyond assigning a uniform terminal reward, we introduce Shortest-Path Reward Adjustment (SPA), which reshapes rewards with respect to task length. The goal of the adjustment is to provide more informative learning signals, guiding the model toward accurate and efficient completion paths and facilitating the computation of trajectory-level advantages.
Second, a further challenge arises from the uniform sampling strategy employed in standard GRPO. In mobile use scenarios—where each sample carries a significant computational cost—this approach results in poor sample efficiency, particularly due to the repeated inclusion of inherently unsolvable tasks. To mitigate this, we adapt data collection and training based on instance difficulty through two mechanisms: Difficulty–Adaptive Positive Replay (AdaPR) and Failure Curriculum Filtering (FCF). At the same time, we restrict redundant successful trajectories to avoid unnecessary updates and promote training efficiency.
2.2.1 Shortest‐Path Reward Adjustment (SPA)
In mobile tasks, the environment returns a binary terminal reward $r\in\{0,1\}$ (Xu et al., 2024; Rawles et al., 2024) indicating task success. Previous approaches typically broadcast this reward to every timestep, i.e., $R(s_{t},a_{t})=r,\;t=0,\dots,T$, so that the per-step signal remains aligned with the sparse objective. However, assigning identical rewards to all successful rollouts biases training toward longer trajectories, since they contribute more gradient terms. To counteract this, we introduce SPA, which re-scales the reward for each trajectory $\tau_{i}$ as
$R^{\mathrm{SPA}}(s_{t},a_{t})=r(\tau_{i})\left(1-\alpha\,\frac{T_{i}-T_{\min}}{T_{i}}\right),\qquad T_{\min}=\min_{\tau_{j}\in\mathcal{T}_{\mathrm{succ}}}|\tau_{j}|,\;\alpha\in(0,1].$ (2)
where $T_{i}=|\tau_{i}|$ is the length of trajectory $\tau_{i}$, and $\mathcal{T}_{\mathrm{succ}}=\{\tau_{j}\mid r(\tau_{j})=1\}$ denotes the set of successful trajectories for the current problem instance. Here $T_{\min}$ is the length of the shortest successful trajectory in $\mathcal{T}_{\mathrm{succ}}$, and $\alpha\!\in(0,1]$ controls the penalty strength. In this formulation, shorter sequences are not automatically considered better; unsuccessful early terminations still receive a reward of 0. This adjustment encourages the policy to prefer shorter, successful paths without sacrificing the success rate.
2.2.2 Difficulty–Adaptive Positive Replay (AdaPR)
In sparse-reward mobile environments, successful yet challenging rollouts are rare but highly informative; effectively leveraging them enhances the learning signal and stabilizes policy updates. Therefore, inspired by the paradigm of experience replay (Mnih et al., 2015; Zha et al., 2019), we introduce difficulty-adaptive positive replay (AdaPR) strategically to retain and reuse challenging, high-value trajectories while blending them with fresh on-policy samples. We introduce key components of AdaPR: buffer construction for high-quality trajectory selection and mixture sampling to balance replay and exploration.
Buffer Construction.
At iteration $t$, the rollout set is $\mathcal{T}_{t}=\{\tau^{(1)},\dots,\tau^{(N)}\}$, collected under the current policy $\pi_{\theta_{t}}$. We compute the trajectory-level advantage and insert the top $\kappa$ trajectories into the replay buffer $\mathcal{B}$.
Mixture Sampling.
Each policy update is performed on a mini‑batch of $M$ trajectories obtained from the mixture distribution $q(\tau)\;=\;\gamma\,p_{\mathcal{B}}(\tau)\;+\;(1-\gamma)\,p_{\mathrm{on}}(\tau)$ where $p_{\mathrm{on}}$ is the on‑policy distribution $\pi_{\theta_{t}}$ and $p_{\mathcal{B}}$ is the empirical distribution over $\mathcal{B}$. To keep the replay contribution under control, at most $\gamma M$ trajectories with the highest current advantage are drawn from $\mathcal{B}$, preserving on‑policy diversity.
2.2.3 Failure Curriculum Filtering (FCF)
To avoid repeatedly sampling tasks that yield zero reward—which wastes computation and hinders the collection of positive advantage data, we propose failure curriculum filtering. In FCF, any task producing all-zero rewards for two consecutive epochs enters a three-epoch cooldown, during which its sampling probability is reduced according to $w_{\text{task}}=\exp(-f)$, where $f$ is the number of consecutive failure epochs; after this cooldown period, the task is permanently removed. This method can be regarded as a variant of curriculum sampling (Matiisen et al., 2019; Narvekar et al., 2020). To avoid excessive hyperparameter tuning, we simplify it to target only the exclusion of the most difficult subset of data. Tasks with consistently high failure counts are permanently removed from the sampling pool. For stability, failure histories from previous training are retained.
Summary.
In summary, MobileRL consists of reasoning-free SFT, reasoning SFT, and difficulty-adaptive RL for training mobile GUI agents. Reasoning-free SFT helps build a strong action foundation from expert demonstrations, while reasoning SFT adds intermediate reasoning to improve instruction following and policy transparency. On top of this initialization, agentic RL with AdaGRPO addresses the challenges of sparse terminal rewards, heavy-tailed task difficulty, and expensive sampling. Specifically, SPA reshapes terminal rewards for denser feedback, AdaPR strategically reuses challenging successful trajectories, and FCF filters out persistently-unsolvable tasks.
3 Experiments
Table 1: Success rates (%) of proprietary and open-source models on AndroidWorld and AndroidLab for mobile GUI interaction tasks.
Models #Params AndroidWorld AndroidLab
Proprietary Models
GPT-4o-2024-11-20 (OpenAI, 2023) - 34.5 31.2
Claude-Sonnet-4-20250514-thinking (Anthropic, 2023) - 41.0 40.6
UI-Tars-1.5 (Qin et al., 2025) - 64.2 38.3
AutoGLM-2024-10 (Liu et al., 2024) - 36.2
Open Models
Qwen2.5-VL-7B-Instruct (Bai et al., 2025) 7B 27.6 10.1
GLM-4.1V-9B-Thinking (Team et al., 2025) 9B 41.7 24.6
UI-Tars-7B (Qin et al., 2025) 7B 33.0 32.6
V-Droid (Dai et al., 2025) 8B 59.5 38.3
UI-Genie-Agent (Xiao et al., 2025) 72B - 41.2
MobileRL (Ours)
w/ Qwen2.5-VL-7B 7B 72.0 42.5
w/ GLM-4.1V-9B-Base 9B 80.2 53.6
3.1 Experiments Settings
Datasets and Benchmarks.
In the two-stage SFT process, we utilize data from human annotations and AndroidControl (Li et al., 2024), constructing 97.9k and 23.6k training steps, respectively. During the RL stage, we employ interactive training sets from AndroidWorld (Rawles et al., 2024) and AndroidLab (Xu et al., 2024), consisting of 2,000 and 1,103 tasks. These were paired with verifiable rewards and VLM-based reward models (cf. Appendix D.3.2). We evaluate on two interactive Mobile benchmarks: AndroidWorld, which includes 116 tasks across 20 apps, and AndroidLab, which comprises 138 tasks across 9 apps. Both benchmarks provide interactive environments. We further assess performance on the static dataset AndroidControl (Li et al., 2024), which contains 8,444 test samples. Due to space limitations, additional details on the training and test datasets are provided in Appendix D.
Baselines.
Our baselines encompass both closed‑ and open‑source agents and models. The closed‑source LLMs include GPT‑4o‑2024‑11‑20 (OpenAI, 2023) and Claude-Sonnet-4-20250514-thinking (Anthropic, 2023), and closed‑source agents UI‑Tars‑1.5 (Qin et al., 2025) and AutoGLM (Liu et al., 2024). The open‑source VLMs, including Qwen2.5‑VL‑7B-Instruct (Bai et al., 2025), GLM-4.1V-9B-Thinking (Team et al., 2025), UI‑Tars‑7B (Qin et al., 2025), V‑Droid (Dai et al., 2025) and UI-Genie-Agent (Xiao et al., 2025), are used.
3.2 Main Results
We evaluate MobileRL by using Qwen2.5-VL-7B and GLM-4.1V-9B-Base as backbones on online interactive benchmarks, AndroidWorld and AndroidLab. As shown in Table 1, MobileRL significantly outperforms previous results. With Qwen2.5-VL-7B as the backbone, MobileRL achieves 72.0%  on AndroidWorld and 42.5%  on AndroidLab, outperforming previous state-of-the-art methods. By using GLM-4.1V-9B as the backbone, the performance of MobileRL further improves to 80.2%  on AndroidWorld and 53.6%  on AndroidLab, achieving the best results across all models.
3.3 Ablation Study
Models AndroidWorld AndroidLab
Qwen2.5-VL-7B-Instruct (Bai et al., 2025) 27.6 10.1
   + Reasoning-Free SFT 50.2+22.6 36.9+26.8
    + Reasoning SFT 56.8+6.6 38.7+1.8
     + AdaGRPO (MobileRL-7B) 72.0+15.2 42.5+3.8
GLM-4.1V-9B-Base (Team et al., 2025) 7.7 10.1
   + Reasoning-Free SFT 48.1+40.4 42.7+32.6
    + Reasoning SFT 66.2+18.1 45.0+2.3
     + AdaGRPO (MobileRL-9B) 80.2+14.0 53.6+8.6
(a) Improvements of task success rates by incrementally applying Reasoning-Free SFT, Reasoning SFT, and AdaGRPO to the base models, respectively.
(b) Training trajectory-level rewards from the AndroidWorld environment with respect to training steps.
Models AndroidWorld
MobileRL 71.1
   w/o AdaPR 63.6
   w/o SPA 69.1
   w/o AdaPR&SPA 58.5
   w/o FCF 64.8
   w/o AdaGRPO 56.8
(c) Test performance on the AndroidWorld test set under different variants.
Figure 3: Ablation studies of the MobileRL framework and its AdaGRPO algorithm. We use the Reasoning SFT model with Qwen2.5-VL-7B-Instruct backbone for the ablation of the AdaGRPO algorithm. All test set results are averaged over three runs to mitigate the impact of randomness.
To evaluate the contributions of the MobileRL framework and the components of the AdaGRPO algorithm, we start with two base models and progressively apply Reasoning-Free SFT, Reasoning SFT, and AdaGRPO. Then, using the Qwen2.5-VL-7B-Instruct model trained with Reasoning SFT as the initialization point and on the AndroidWorld training set, we conduct an analysis of the impact of each component of AdaGRPO—AdaPR, SPA, and FCF.
MobileRL Ablation.
We summarize stage-wise gains in success rate (SR) in Table 3(a). For Qwen2.5-VL-7B, the combined improvements of MobileRL are +44.4% on AndroidWorld and +32.4% on AndroidLab; For GLM-4.1V-9B, the overall gains are +72.5% and +43.5%. Overall, Reasoning-Free SFT delivers a strong initial lift, and Reasoning SFT offers additional improvements. Building upon the strong foundation established by the preceding stages, the AdaGRPO stage further augments the final performance, achieving an improvement exceeding 10% on the AndroidWorld dataset and over 5% on the AndroidLab test set.
AdaGRPO Ablation.
The design of AdaGRPO covers SPA, AdaPR, and FCF. The ablations are performed on four settings: (i) MobileRL w/o AdaPR (no replay), (ii) MobileRL w/o SPA (no reward shaping), (iii) MobileRL w/o AdaPR & SPA (neither), and (iv) MobileRL w/o FCF (uniform sampling).
We report on-policy trajectory reward curves during training (excluding replayed trajectories) in Figure 3(b) and the success rates on the AndroidWorld in Table 3(c). To avoid bias from the AndroidLab reward model, we use only AndroidWorld in this study. Each component of AdaGRPO contributes to improving the performance of MobileRL. Specifically, we have the following observations:
  • FCF under constraints. With a 100-step budget (>40>40 hours), FCF plays a key role in filtering. Removing it biases early sampling toward overly hard tasks, yielding many negatives and a lower reward ceiling. The w/o FCF curve keeps rising, suggesting stronger results with more efficient pipelines or simulators.
  • FCF only (w/o AdaPR & SPA). Training is initially stable but collapses after about 30 steps, indicating that AdaPR and SPA are necessary for stabilizing the training.
  • Effect of AdaPR. After about seven steps (once the replay buffer is populated), the gap between w/o AdaPR and the full MobileRL method grows, showing the benefit of replay.
  • Effect of SPA. Noticeable gains of SPA appear after roughly 60 steps, likely because the lack of step-wise control leads to overly long trajectories.
Is Reasoning-Free SFT Still Necessary?
We apply supervised fine-tuning on the expert dataset without reasoning traces, which we term Reasoning-Free SFT. This raises a key question: Is fine-tuning with expert data that lacks reasoning still beneficial?
We compare MobileRL (Reasoning-Free SFT + Reasoning SFT + AdaGRPO) with MobileRL without Reasoning-Free SFT in Figure 3. Interestingly, our experiments show that incorporating Reasoning-Free SFT consistently improves performance on AndroidWorld. It suggests that without explicit reasoning traces, (Reasoning-Free) SFT contributes to enhancing final results for MobileRL.
Table 2: Results on the AndroidControl (Li et al., 2024). Note that data from AndroidControl was not included during the RL stage.
Model AndroidControl-Low AndroidControl-High
Type Grounding SR Type Grounding SR
UI-Tars-7B 98.0 89.3 90.8 83.7 80.5 72.5
UI-Tars-72B 98.1 90.8 91.3 85.2 81.5 74.7
UI-Genie-7B 98.1 89.3 94.3 83.5 82.9 74.5
UI-Genie-72B 98.3 95.4 94.8 84.9 86.3 77.0
MobileRL-7B
   w/o RL 96.7 93.7 91.9 87.0 73.6 72.8
   w/ RL 96.9 92.4 91.4 86.3 71.5 71.3
MobileRL-9B
   w/o RL 98.6 95.8 95.9 89.8 78.1 77.5
   w/ RL 97.9 94.6 94.3 87.0 74.5 73.9
Table 3: Effect of reasoning-free SFT evaluated on AndroidWorld.
Performance on Offline Dataset.
As shown in Appendix D, Reasoning-Free and Reasoning SFT stages include training data from AndroidControl, which has been converted into MobileRL format. In Table 3, we present the results of MobileRL on the AndroidControl test set. The SR score of MobileRL-9B w/o RL version surpasses all previous models, achieving the state of the art, while the RL version largely maintains the score.
Success Rates by Task Complexity.
We divide the AndroidWorld test set by rounded-up Complexity (Rawles et al., 2024): complexity level 1 ($\leq$10 steps), level 2 (11–20), level 3 (21–30), and level 4+ (>>30). We run eight test trials at temperature $1.0$ and report pass@1/2/4/8 in Figure 5. MobileRL yields consistent gains at all complexity levels. Notably, post-RL pass@1 exceeds pre-RL pass@8, indicating substantial effectiveness. Consistent with AdaPR’s design for heavy-tailed difficulty, larger gains occur on high-complexity tasks.
Figure 4: Pass@$k$ on AndroidWorld by task complexity (Rawles et al. (2024)) levels. Pass@$k$ is the fraction of tasks solved within the top-$k$ attempts.
Figure 5: Win rate of MobileRL vs. MobileRL w/o SPA, where a win means completing a task with fewer steps. $n$ denotes the number of task templates per category. Categories: All (all templates); C1–C4 (complexity levels 1–4); BC/BW (both methods correct/wrong); Others (exactly one method correct).
Impact of SPA on Step Efficiency.
Although SPA has the smallest impact on overall accuracy in the ablation, its effect on step efficiency is clear. As shown in Figure 5, partitioning tasks by complexity reveals that MobileRL with SPA consistently completes tasks in fewer steps across all difficulty levels. Moreover, when we compare cases where both models are correct (BC), both are wrong (BW), and those where only one is correct, MobileRL with SPA more frequently yields shorter trajectories in every group.
4 Related Work
Mobile GUI Agents.
Recent work leverages powerful language models to build agents that operate real PCs and phones (Agashe et al., 2025; Qin et al., 2025; Lai et al., 2025b), including Android agents that perceive GUIs and act via taps, swipes, and text (Toyama et al., 2021; Xu et al., 2024). To improve action prediction and learning, frameworks explore multimodal exploration (Yang et al., 2023), modular reasoning (Lai et al., 2025a), verifier-driven control (Dai et al., 2025), and small-LM code-based execution (Wen et al., 2025). Yet many systems still rely on offline RL or single-turn data: DigiRL uses offline demonstrations (Bai et al., 2024); U1-R1 trains on single-step episodes (Lu et al., 2025); and UI-Tars applies DPO in an offline regime (Qin et al., 2025), leaving online, multi-turn RL for adaptive mobile agents unexplored. Studies emphasize realistic applications: AppAgent (Yang et al., 2023) evaluates closed-source models on real-world apps, while A3 (Chai et al., 2025) offers a realistic app suite with an autonomous evaluation protocol that reduces human effort.
Benchmarks for Mobile Agents.
Benchmarking generally follows two tracks. Static or replay-based settings—AndroidControl (Li et al., 2024), Android in the Wild (Rawles et al., 2023), MobileAgentBench (Wang et al., 2024), and Mobile-Bench (Deng et al., 2024) offer plenty of tasks and trajectories. These benchmarks fall short for real-world evaluation, as fixed trajectories and screens limit agents’ ability to handle uncertainty and exploration. Interactive emulator environments—AndroidWorld (Rawles et al., 2024), AndroidLab (Xu et al., 2024), and B-MOCA (Lee et al., 2024)—span diverse apps and realistic tasks, yet remain challenging for current agents. Current mobile GUI benchmarks are limited: they lack asynchronous, parallel VM interaction for scalable training. To the best of our knowledge, all public mobile GUI benchmarks to date target the Android operating system.
5 Conclusion
In this work, we present MobileRL, an agentic RL framework that advances mobile GUI agents. It achieves this by combining staged initialization with an adaptive reinforcement learning algorithm (AdaGRPO). Training begins with reasoning-free SFT on large-scale expert demonstrations, followed by a reasoning SFT stage that adds intermediate rationales and reduces cold-start exploration costs. Building on this, we introduce Difficulty–ADAptive GRPO (AdaGRPO), which enhances GRPO with shortest-path reward adjustment, adaptive positive replay, and failure curriculum filtering. These three strategies improve sample efficiency and guide policies toward more accurate and efficient task completion. Experiments on AndroidWorld and AndroidLab demonstrate that MobileRL with open models significantly outperforms both open-source and proprietary baselines.
Appendix A Case Study
(a) Add a location marker for 47.16, 9.51 in the OsmAnd maps app.
(b) Search for wireless headphones in temu, and sort by price low to high.
(c) Search for hotels on Booking, check-in date is 09-10, check-out date is 09-12, sort by prices.
Figure 6: Example mobile tasks finished by our agent. Our agent can automatically perform tasks according to human instructions in academic benchmarks and real-world applications.
Figure 7: The Reasoning-Free SFT agent, where the expert data does not include the reasoning part, in the SimpleCalendarAddRepeatingEvent task in AndroidWorld. Because of space constraints, we only present the most critical and notable steps for comparison.
Figure 8: Reasoning SFT agent in the SimpleCalendarAddRepeatingEvent task in AndroidWorld. Because of space constraints, we only present the most critical and notable steps for comparison.
Figure 9: MobileRL in the SimpleCalendarAddRepeatingEvent task in AndroidWorld. Because of space constraints, we only present the most critical and notable steps for comparison.
We present a case study from AndroidWorld evaluating three agents: the Reasoning-Free SFT agent, Reasoning SFT agent, and MobileRL agent. We choose SimpleCalendarAddRepeatingEvent, which requires creating a recurring event titled “Review session for Budget Planning,” starting on 2023-10-15 at 14:00, lasting 60 minutes, repeating weekly without end, and including the description: “We will understand business objectives. Remember to confirm attendance.”
All agents successfully configured the title and basic settings in the initial steps. Their performance diverged in subsequent steps. The Reasoning-Free SFT agent made timing errors (setting 16:00 instead of 15:00) and executed redundant checks, revealing weak task understanding (Figure 7). The Reasoning SFT agent skipped an adjustment step, yielding an incorrect event duration (Figure 8). By contrast, the full MobileRL agent completed the task accurately and efficiently, satisfying all requirements without redundant operations (Figure 9).
Appendix B Training Details
B.1 Hyperparameters
In both the Reasoning-Free SFT and Reasoning SFT stages, we conducted fine-tuning for two epochs, employing a cosine-decayed learning rate scheduled from $1\times 10^{-5}$ to $1\times 10^{-6}$. The training was implemented using Llama-Factory, with packing mode enabled to accelerate training efficiency. All images were used at their full resolution without compression.
In the RL stage, we extend the GRPO framework from Verl with our customized functionalities. The main hyperparameters are summarized in Table 4.
Table 4: Summary of Main Hyperparameters
Component Hyperparameter Value
Data Max Prompt Length 16384
Data Max Response Length 4096
Data Train Batch Size 256
Data Validation Batch Size 256
Actor / Policy Strategy (Parallelism) FSDP
Actor / Policy PPO Micro Batch Size/GPU 4
Actor / Policy Learning Rate (LR) 1e-6
Actor / Policy Gradient Clipping 1.0
Actor / Policy Clip Ratio 0.2
Actor / Policy PPO Epochs 1
Rollout & Sampling Sampling Temperature 1.0
Rollout & Sampling Max New Tokens 4096
Rollout & Sampling Number of Samples (n) 16
Rollout & Sampling Max Turns 50
Rollout & Sampling Max Pixels 5000000
Rollout & Sampling Min Pixels 65536
Algorithm KL Loss Coefficient/ $\beta$ 0.001
Algorithm SPA/ $\alpha$ 1.0
Algorithm AdaPR/ Replay Buffer Size 256
Algorithm AdaPR/ $\gamma$ 1.0
Algorithm AdaPR/ $\kappa$ 0.25
Table 5: Action Space for Mobile GUI Interaction. We utilize a merging action space from AndroidLab (Xu et al., 2024) and AndroidControl Li et al. (2024), which represents screen positions with bounding boxes aligned to XML data.
Action Parameters Description
Tap element=[x1,y1,x2,y2] Tap at the rectangle defined by top‐left (x1,y1) and bottom‐right (x2,y2).
Type text={string} Enter the given string into the focused input field.
Swipe
direction={up/down/left/right}
dist={short/medium/long}
element=[x1,y1,x2,y2] (optional)
Swipe in the given direction over the specified distance. Optionally constrain to the rectangle element.
Long Press element=[x1,y1,x2,y2] Press and hold on the given rectangle area.
Launch app={AppName} Launch the named application.
Back none Press the system Back button.
Home none Press the system Home button.
Wait none Wait for three seconds.
Finish message={string} (optional) End the session with an optional message.
B.2 Action Space.
We design a series of actions based on AndroidLab (Xu et al., 2024), which supports a concise set of actions for GUI interaction, as described in Table 5.
B.3 Implement Details
Pruning Negative Trajectories
To stabilize training, we prune trajectories with the lowest advantages, reducing noisy samples in the replay buffer. High-advantage trajectories, if stored, are expected to be sampled only once on average. Thus, we maintain a maximum positive-to-negative trajectory ratio of 1:2 by randomly discarding the lowest-advantage trajectories.
We further study a pruning strategy that discards overly frequent erroneous trajectories from the reinforcement learning buffer before each update. As depicted in Figure 10, this filtering keeps the policy entropy consistently higher during training. By pruning trajectories whose advantages remain persistently negative, the agent avoids being driven by detrimental gradients; probability mass is instead spread over a broader action space, fostering exploration and delaying premature convergence. This pruning strategy ultimately leads to more robust policy learning.
Figure 10: Effect of the pruning-negative strategy evaluated on AndroidWorld.
Influence of Image Resolution
Table 6: Comparison of success rate between compressed and original images across different training stages and datasets.
Model Compressed Images Original Images
AndroidWorld AndroidLab AndroidWorld AndroidLab
Reasoning-Free SFT 48.9 39.8 48.1 42.7
Reasoning SFT 66.1 40.3 66.2 45.0
AdaGRPO (MobileRL-9B) 75.8 46.8 80.2 53.6
In our preliminary experiments, we applied a uniform compression to images (maximum pixels per image = 500,000) across all three stages in order to accelerate training, yielding the initial results. The standard practice, however, is to use uncompressed images. For experiments based on Qwen2.5-VL-Instruct, we observed that this difference did not significantly affect performance while greatly accelerating training. In contrast, for experiments based on GLM-4.1V-9B-Base, we found notable accuracy differences between compressed and uncompressed settings, as shown in Table 6. Therefore, we updated the reported results to those obtained with full-resolution images.
This change in accuracy is reasonable in the SFT stage, since the AndroidLab tasks contain more question-answering oriented tasks, where higher resolution images often make it easier to read screen content. In the RL stage, clearer images allow the agent to explore more thoroughly, leading to higher scores.
Appendix C XML Preprocessing for UI Representation
The original XML from the Android accessibility service defines the layout and elements of the user interface, including all components on a page. As a result, it contains many nodes used solely for structural or layout purposes, which do not provide useful semantic information. Moreover, scrollable pages often contain more nodes than are visible on the screen, leading to the inclusion of many off-screen nodes.
C.1 Removal of Off-Screen Nodes
We first determine whether to retain off-screen nodes via the input parameter remain_nodes:
  • remain_nodes=True: Off-screen nodes are preserved, e.g., when summarizing the full page content without requiring scrolling.
  • remain_nodes=False: Off-screen nodes are removed to avoid interference during action simulation (e.g., tapping, scrolling).
In the original XML, a node is considered on-screen if its bounds property lies entirely within the screen dimensions $[0,0]$ to [Window_Height,Window_Width][Window\_Height,Window\_Width] and is contained by its parent node. We check this condition recursively to identify on-screen nodes.
C.2 Removal of Redundant Nodes
Nodes that do not convey functional or semantic information are removed. A node is considered functional if it satisfies at least one of the following:
  • Any of the boolean attributes is True: checkable, checked, clickable, focusable, scrollable, long-clickable, password, selected.
  • The text or content-desc attribute is non-empty.
All nodes failing these criteria are considered redundant and are deleted.
C.3 Attribute Simplification
Attribute descriptions in the original XML are verbose and token-expensive. We simplify them as follows:
  • Keep only True values for the boolean functional attributes listed above (omit False values).
  • Remove index, resource-id, and package (not useful for semantic understanding).
  • For class, retain only the last component (e.g., android.widget.FrameLayout $\rightarrow$ FrameLayout).
  • Merge text and content-desc attributes and display them separately.
  • Retain bounds in full, as it indicates the node’s position on the page.
C.4 Example Transformation
Original node:
<node index="0" text="Audio Recorder"
resource-id="com.dimowner.audiorecorder:id/txt_title"
class="android.widget.TextView" package="com.dimowner.audiorecorder"
content-desc="" checkable="false" checked="false" clickable="false"
enabled="true" focusable="false" focused="false" scrollable="false"
long-clickable="false" password="false" selected="false"
bounds="[221,1095][858,1222]" />
Simplified node:
TextView;;Audio Recorder;[221,1095][858,1222]
Appendix D Data Collection
In this section, we present the composition and construction methodology of the training and testing data used at each stage, as well as their respective proportions.
D.1 Data Distribution
We present the source distribution of the first two stages of SFT data in Table 7. In the RL stage, we utilized 2,000 tasks from AndroidWorld and 1,103 tasks from AndroidLab. Details of our data construction, deduplication, and reward acquisition methods are provided in a later section. We compared the performance of training solely on AndroidWorld with that of mixed training on both AndroidWorld and AndroidLab. The results varied across different model backbones: models based on Qwen2.5-VL-7B-Instruct performed better under the mixed-training setting, whereas models based on GLM-4.1V-9B-Base achieved superior results when trained only on AndroidWorld. Accordingly, we report the best results for each backbone.
Table 7: Data distribution and labels across the two SFT training stages by steps.
Stage
Android
Control (Low)
Android
Control (High)
Human Annotation
Total
Reasoning-Free SFT 21.3k 14.3k 62.2k 97.9k
Reasoning SFT 7.2k 4.1k 12.2k 23.6k
D.2 Human Annotation
Following the approach in Xu et al. (2024), we collected a portion of human-annotation data and additionally employed model-driven self-exploration on the online interactive environment, performing online rollouts and selecting the correct action sequences for training.
Privacy Protection
In this work, the data collection process for model training inevitably involves capturing page screenshots and corresponding XML files. To address this, a privacy protection interface that does not retain raw data is applied before storage, ensuring that all sensitive information is filtered out in advance and that only sanitized data is preserved. Building upon this foundation, the overall scope of privacy protection is defined to cover both textual and visual information, with sensitive elements being systematically identified and anonymized. For textual data, the detection process primarily employs regular expression–based rules to recognize personal identifiers such as social media accounts, vehicle registration numbers, shipment tracking codes, order numbers, bank account details, email addresses, passport information, national identification numbers, and phone numbers. Once detected, these elements are replaced with standardized masking tokens that indicate the category of the information without revealing the actual content, thereby preserving semantic integrity while mitigating risks of exposure. For visual data, the privacy protection mechanism extends to the automatic detection of faces, QR codes, and text extracted from images through optical character recognition. Detected sensitive regions within images are then masked or obfuscated, producing new visual outputs in which private content is effectively concealed. Taken together, these measures form a comprehensive privacy-preservation strategy that integrates rule-based textual analysis with multimodal image processing, ensuring robust safeguards against the leakage of personal or identifying information throughout the data collection and annotation pipeline.
D.3 AndroidLab
D.3.1 Data Processing
The AndroidLab (Xu et al., 2024) dataset consists of an online evaluation set containing 138 problems, all of which are equipped with verifiable evaluation mechanisms. In our study, we directly adopt the SR (Success Rate) metric provided by AndroidLab as the primary reward indicator. However, since AndroidLab does not include a corresponding verifiable training set, we construct one by means of a model-driven expansion strategy. Specifically, we utilize the applications included in AndroidLab to automatically generate candidate training tasks, which are subsequently subjected to manual verification to ensure both feasibility and non-overlap with the test set. Through this process, we compile a total of 1,103 training problems.
Since ground-truth test methods do not accompany these generated problems, it is not possible to rely on the same direct evaluation mechanism used for the test set. To address this issue, we train a reward model that serves as an auxiliary evaluator of correctness. This reward model is employed in two stages of our training pipeline: (i) during the Reasoning SFT stage, it functions as a filter to ensure the quality and validity of training data; and (ii) during the Reinforcement Learning stage, it provides a learnable reward signal that guides policy optimization.
D.3.2 Reward Model for AndroidLab
Since the AndroidLab environment does not provide rule-based rewards for training data, we adopt a VLM-based reward model to supply reward signals during reinforcement learning. Specifically, we first execute all training and test tasks multiple times using medium-capability VLMs, including different versions of our models and proprietary VLMs, generating execution traces. Then, strong proprietary VLMs assign binary scores to each trace. For scoring, each model receives the full task screenshot sequence concatenated into a single image, along with step-by-step action descriptions, and is instructed to produce a reasoning process before outputting a score. For training-set tasks (non-rule-based), we use the majority vote among the three scores as the label and retain the reasoning; for test-set tasks, we use the existing rule-based reward as the label and keep only the model scores and reasoning consistent with it. This process yields several thousand labeled samples, which we use to fine-tune GLM-4.1V-9B-Thinking as the base model. After training, we evaluate the reward model on a curated set of 1000 AndroidLab traces with verified labels, selecting the best-performing version (86% accuracy) for online reinforcement learning.
System Prompt for Trace Evaluation
The following system prompt guides the VLM in determining whether an agent has successfully completed a task.
You are an expert in determining whether a task
has been successfully and completely completed. You will receive:
1. The task description.
2. Step-by-step page states in XML format.
3. The agents action descriptions.
4. A single image containing screenshots of all steps.
- Green rectangles and crosses mark tap regions and positions.
- Green arrows indicate swipe directions.
- Red text shows typed input.
Action formats and meanings:
- do(action="Tap", element=[x1,y1,x2,y2])
Tap on the specified screen region; green cross is the tap point.
- do(action="Launch", app="xxx")
Launch the specified app.
- do(action="Type", text="xxx")
Enter the specified text (shown in red).
- do(action="Swipe", element=[x1,y1,x2,y2], direction="x", dist="x")
Swipe in the indicated direction; green arrow shows the swipe path.
- do(action="Long_Pres", element=[x1,y1,x2,y2])
Long press on the specified region.
- do(action="Back")
Navigate back to the previous screen.
- finish(message="xxx")
End the task with the given message.
Scoring rules:
If the task is fully and correctly completed,
output a score of 1; otherwise, 0.
Output format:
<analysis>
Step 1 analysis: <Your analysis>
Step 2 analysis: <Your analysis>
Final step analysis: <Your analysis>
</analysis>
<ans>
[Your score]
</ans>
Analysis
Figure 11: Trajectory-level rewards with 95% CIs on training sets, with the reasoning SFT version of Qwen-2.5-VL-Instruct as backbone model. showing consistent performance growth.
We compared the training curves of AndroidWorld and AndroidLab (trained simultaneously), as shown in Figure 11. The training curve of AndroidWorld is smoother compared to that of AndroidLab, which we attribute to the use of a rule-based reward in AndroidWorld, whereas AndroidLab relies on a reward model. However, the reward model cannot fully match the accuracy of the rule-based reward. Therefore, we believe that evaluating whether a broader range of tasks has been successfully completed without relying on rule-based rewards remains an important challenge for future work.
D.4 AndroidWorld
D.4.1 Data Processing
The AndroidWorld dataset (Rawles et al., 2024) comprises 116 test problems, each equipped with verifiable reward signals. Since its native action space differs from ours, we adapt the operation code to align with our action space. AndroidWorld further provides randomized initialization parameters, which enable the generation of a large number of training tasks with verifiable rewards by replacing predefined substitutable components in the tasks (e.g., input content) and by varying the initial device states. To prevent overlap with the evaluation set, we remove all tasks whose specific content and initial states coincide with those in the test set. From the remaining pool, we curate 2,000 tasks to serve as the training set.
D.5 AndroidControl
Table 8: The correspondence between the AndroidControl action space and MobileRL action space, as well as the judgment details for each action.
AC Action MobileRL Action Evaluation Details
click Tap Consider correct if the action type matches and the click point falls within the Tap bounding box.
long_press Long Press Consider correct if the action type matches and the long_press point falls within the Long Press bounding box.
scroll Swipe Consider correct if the action type matches and the scroll direction corresponds to the Swipe direction.
open_app Launch Consider correct if the action type matches and the Launch app matches.
input_text Type Consider correct if the action type matches and the Type text F1 score with the input_text text is higher than 0.5.
navigate_home Home Consider correct if the action type matches.
navigate_back Back Consider correct if the action type matches.
wait Wait Consider correct if the action type matches.
D.5.1 Data Processing
Each test case in the AndroidControl (Li et al., 2024) dataset includes episode_id, goal, screenshots, accessibility_trees, screenshot_widths, actions, screenshot_heights, and step_instructions. The format of input information and output actions differs from that of our action space, and we provide transformation details in Table 8. Certain transformations are required for an evaluation to be performed. The dataset consists of 8,444 test samples, 690 validation samples, and 74,714 training samples.
Transformation from Accessibility Tree to XML
We first transform the input information accessibility_tree to XML. The XML obtained from mobile pages contains nodes with the following attributes: index, resource-id, class, package, content-desc, checkable, checked, clickable, enabled, focusable, focused, scrollable, long-clickable, password, selected, and bounds. We extract and set these corresponding attributes from each node of the accessibility tree. This ensures that the converted XML is consistent with the XML obtained directly from mobile pages. We recursively process through steps as follows:
  • Recursively process the three levels of accessibility_tree, window, and node.
  • Set the corresponding XML attribute for each attribute of the node.
  • Compress the converted XML into compressed_XML format.
Here is a transformation example:
Accessibility tree node:
nodes {
bounds_in_screen {
right: 1080
bottom: 2400
class_name: "android.widget.FrameLayout"
package_name: "com.zoho.meeting"
text_selection_start: -1
text_selection_end: -1
window_id: 1733
is_enabled: true
is_visible_to_user: true
actions {
id: 4
actions {
id: 8
actions {
id: 64
actions {
id: 16908342
XML node:
<node index="0" text="" resource-id="" class="android.widget.FrameLayout" package="com.zoho.meeting" content-desc="" checkable="false" checked="false" clickable="false" enabled="true" focusable="false" focused="false" scrollable="false" long-clickable="false" password="false" selected="false" bounds="[0,0][1080,2400]"/>
Transformation from Point to Bounding Box
The grounding actions in AndroidControl use points. Since our grounding actions use bounding boxes rather than points, conversion of grounding coordinates is required. We attempt to match all UI boxes on a page. If the point falls within a UI box, the original click coordinates are converted to the coordinates of that UI box.
Evaluation Details
Our evaluation metric follows Lian et al. (2025), as it provides open-source complete evaluation code, as shown in Table 8. Notably, since our grounding actions use bounding boxes instead of points, we have modified the judgment logic for positioning operations. A grounding action is considered correct if the point coordinates of the original action fall within the UI box predicted by the model. For the AndroidControl Low evaluation, we only conduct single-turn conversations with the agent. The provided information includes the instruction for the current step, the screenshot, and compressed XML data. For the AndroidControl High evaluation, we maintain a history with the agent and replace incorrect actions. The provided information includes the total task, the screenshot for the current step, and compressed XML data.
Appendix E Limitation
Reward Model
We train a reward model to gain reward signals during the training on the AndroidLab environment. The training data is obtained from strong proprietary VLMs through majority voting, but it is noteworthy that bias still exists in those data. Therefore, the reward model is inevitably born slightly biased from a perfect judge on agentic tasks, which brings about unsatisfying results during online RL in some cases. In particular, while MobileRL-7B and all ablation studies are carried out with the existence of the reward model, the best model, MobileRL-9B, is gained from online RL purely on the AndroidWorld environment. In observation of this, we conclude that a general and unbiased verification model is of significant importance and necessity, which is proposed as one aspect of our future work.
Difficulties with Public-Use Apps
Our current work does not evaluate on out-of-domain public-use apps. Although the training data from AndroidControl and human annotations both include some tasks within public-use apps, we did not conduct testing in this setting. The main reason is the lack of a reproducible and efficient evaluation protocol in the open domain. Existing approaches typically require human execution of tasks and manual assessment of outcomes, which are costly and difficult to scale.
Moreover, public-use apps often introduce practical challenges. Many require user login, cannot be run reliably on virtual machines, or lack mechanisms to restore application states, making online rollouts particularly challenging from an engineering perspective. We plan to explore solutions for training and evaluating on public-use apps in future work.
Limitation of prediction based on XML
We observed that certain applications on the Android platform, especially those with large popularity and of comprehensive functions(e.g. TikTok, etc.), lack sufficient information about all interactable elements in their XML hierarchy data. To successfully perform complex tasks on these Apps, the current format we chose in model prediction, which is based on bounding box information contained in XML data, needs adaptation toward the general format built on solely visual information. This adaptation is, in the end, inevitable while challenging, so we leave it for our future work to obtain a generalized vision model with no need for XML data.
Training Time and Resource Consumption
Another limitation of our work lies in the substantial resource demands of Android virtual devices. Each device requires more than 5 GB of memory and over 30 GB of storage, which restricts the number of rollouts that can be run concurrently. To address this, we employed four CPU machines with 1 TB of memory each, dedicated solely to rollouts. Although we implemented a distributed framework that supports multi-machine concurrent sampling, the system can only stably sustain up to 256 parallel rollouts, with each step taking more than 10 minutes.
This constraint severely limits further scaling of training steps, as training 100 steps already takes more than 25 hours. In addition, large-scale image inference and network transmission introduce significant overhead, further increasing the computational burden.